# GenMol: A Drug Discovery Generalist with Discrete Diffusion

### v1

@@ 4.2 Exploration with Fragment Remasking
Under the discrete diffusion framework of GenMol, we propose fragment remasking (Figure 3), an effective strategy to explore the neighborhood of a given molecule in chemical space and find optimized molecules. Discrete diffusion allows GenMol to mask and re-predict some tokens of a given molecule, making neighborhood exploration simple and straightforward. However, although there exist some works that apply token-wise remasking to protein sequence optimization (Hayes et al., 2024; Gruver et al., 2024), it would be suboptimal to naively adopt the same strategy for small molecule optimization. This is because each token in a SAFE (or SMILES) sequence represents a single atom or bond, and masking them individually results in localized and ineffective exploration. The structure and activity of molecules are highly correlated, which is known as the structure-activity relationship (SAR) (Bohacek et al., 1996), and the units that carry information about molecular properties are fragments, not individual atoms or bonds. Inspired by the principle of SAR, fragment remasking modifies molecules by randomly selecting a fragment in a given molecule and replacing it with a fragment mask chunk, from which discrete diffusion generates a new molecular fragment. The proposed fragment remasking can also be viewed as a mutation operation in GA where the mutation is performed at the fragment-level rather than at the atom- or bond-level. We will examine the superiority of fragment remasking over token remasking in Section 5.5.
@@
The lengths of the fragment mask chunks are sampled from a predefined distribution, e.g., the distribution of fragment lengths in the training molecule. The length of the fragment depends on the decomposition rule used, and instead of using a fixed length, this strategy allows GenMol to automatically adjust the length based on the decomposition rule. This also ensures that GenMol generates fragments of varying lengths, offering users better controllability over molecule generation.
5 Experiments
GenMol is trained on the SAFE dataset (Noutahi et al., 2024), which combines molecules from ZINC (Irwin & Shoichet, 2005) and UniChem (Chambers et al., 2013). With this single trained model, we demonstrate that GenMol can be universally and effectively applied to a wide range of molecule generation tasks simulating various real-world drug discovery scenarios. We emphasize that a single GenMol checkpoint is used to perform the following tasks without any additional finetuning specific to each task. We first conduct experiments on a de novo molecule generation task in Section 5.1. Next, we conduct experiments on fragment-constrained molecule generation tasks in Section 5.2. We then examine GenMol’s ability to perform goal-directed hit generation and goal-directed lead optimization in Section 5.3 and Section 5.4, respectively. We perform ablation studies in Section 5.5.

### v2

@@ 4.2 Exploration in Chemical Space with GenMol
Therefore, utilizing discrete diffusion of GenMol, we propose fragment remasking, an effective strategy to explore the neighborhood of a given molecule in chemical space to find optimal molecules. Discrete diffusion allows GenMol to mask and re-predict some tokens of a given molecule, making neighborhood exploration simple and straightforward. However, although there exist some works that apply token-wise remasking to protein sequence optimization (Hayes et al., 2024; Gruver et al., 2024), it would be suboptimal to naively adopt the same strategy for small molecule optimization. This is because each token in a SAFE (or SMILES) sequence represents a single atom or bond, and masking them individually results in localized and ineffective exploration. The units that carry information about molecular properties are fragments or functional groups, not individual atoms or bonds (Li, 2020), and thus fragment remasking modifies a molecule by randomly re-predicting a fragment.
Given $\boldsymbol{x}_{\text{init}}$, $\boldsymbol{x}_{\text{mask}}$ is constructed by randomly selecting one of the fragments of $\boldsymbol{x}_{\text{init}}$ and replacing it with a mask chunk. Here, a finer decomposition rule than the one that constructed the vocabulary is used, allowing fragment remasking to operate at a fine-grained level. GenMol then generates $\boldsymbol{x}_{\text{new}}$ by iteratively unmasking $\boldsymbol{x}_{\text{mask}}$. The proposed fragment remasking can also be viewed as a mutation operation in GA where the mutation is performed at the fragment-level rather than at the atom- or bond-level. $\boldsymbol{x}_{\text{new}}$ is decomposed and scored by Eq. (5), and the fragment vocabulary is dynamically updated by selecting the top-$V$ fragments, allowing exploration beyond the initial fragments. We summarize the goal-directed generation process of GenMol with fragment remasking in Algorithm 1.
Algorithm 1 Goal-directed Molecular Optimization of GenMol
  Input: A set of molecules $\mathcal{D}$, vocabulary size $V$,    decomposition rule for fragment vocabulary $R_{\text{vocab}}$,    decomposition rule for fragment remasking $R_{\text{remask}}$,    number of generations $G$
  Set $\mathcal{F}\leftarrow$ fragments obtained by decomposing $\mathcal{D}$ with $R_{\text{vocab}}$
  Set $\mathcal{V}\leftarrow$ top-$V$ fragments of $\mathcal{F}$ (Eq. 5)
  Set $p_{\text{len}}\leftarrow$ fragment length distribution of $\mathcal{D}$ based on $R_{\text{remask}}$
  Set $\mathcal{M}\leftarrow\emptyset$
  while $|\mathcal{M}|<G$ do
     Select and attach two fragments from $\mathcal{V}$ to get $\boldsymbol{x}_{\text{init}}$
     Sample the fragment length $m\sim p_{\text{len}}$
     Select one of the fragments of $\boldsymbol{x}_{\text{init}}$ based on $R_{\text{remask}}$ and  replace it with $m$ mask tokens to get $\boldsymbol{x}_{\text{mask}}$
     Generate $\boldsymbol{x}_{\text{new}}$ by iteratively unmasking $\boldsymbol{x}_{\text{mask}}$ with GenMol
     Update $\mathcal{M}\leftarrow\mathcal{M}\cup\{\boldsymbol{x}_{\text{new}}\}$
     Decompose $\boldsymbol{x}_{\text{new}}$ into fragments $\{\boldsymbol{f}_{1},\boldsymbol{f}_{2},\dots\}$ with $\mathcal{R}_{\text{vocab}}$
     Update $\mathcal{V}\leftarrow$ top-$V$ fragments from $\mathcal{V}\cup\{\boldsymbol{f}_{1},\boldsymbol{f}_{2},\dots\}$
  end while
  Output: Generated molecules $\mathcal{M}$
@@
Instead of using a fixed length, the lengths of the fragment mask chunks are sampled from a predefined distribution, e.g., the distribution of fragment lengths in the training set. The length of the fragment depends on the decomposition rule used, and this strategy allows GenMol to automatically adjust the length based on the rule. This also ensures that GenMol generates fragments of varying lengths, offering users better controllability over molecule generation.
Fragment remasking can also be interpreted as Gibbs sampling (Geman & Geman, 1984). Assuming a SAFE molecular sequence $\boldsymbol{x}$ is comprised of $F$ fragments, we can represent $\boldsymbol{x}$ as a set of the fragments $\{\boldsymbol{f}_{k}\}^{F}_{k=1}$, where $\boldsymbol{f}_{k}$ denotes an attachment point-assigned SAFE fragment. To sample a molecule $\boldsymbol{x}$ from $p(\boldsymbol{x})=p(\boldsymbol{f}_{1},\dots,\boldsymbol{f}_{F})$, fragment remasking repeats the process of uniformly selecting the index $k$ and then sampling $\boldsymbol{f}_{k}$ from $p(\boldsymbol{f}_{k}|\boldsymbol{f}_{\backslash k})$, where $\boldsymbol{f}_{\backslash k}$ denotes $\boldsymbol{f}_{1},\dots,\boldsymbol{f}_{F}$ but with $\boldsymbol{f}_{k}$ omitted. This is equivalent to performing Gibbs sampling with the Markov kernel $p(\boldsymbol{f}_{k}|\boldsymbol{f}_{\backslash k})$, allowing GenMol to perform a random walk in the neighborhood of the given molecule $\boldsymbol{x}$.
@@
4.3 Molecular Context Guidance
Inspired by autoguidance (Karras et al., 2024), we propose molecular context guidance (MCG), a guidance method tailored for masked discrete diffusion of GenMol. Karras et al. (2024) generalized classifier-free guidance (CFG) (Ho & Salimans, 2021) and proposed to extrapolate between the predictions of two denoising networks:
$D^{(w)}(\boldsymbol{z}_{s}|\boldsymbol{z}_{t},\boldsymbol{y})=wD_{1}(\boldsymbol{z}_{s}|\boldsymbol{z}_{t},\boldsymbol{y})+(1-w)D_{0}(\boldsymbol{z}_{s}|\boldsymbol{z}_{t},\boldsymbol{y}),$ (6)
where $D_{1}$ and $D_{0}$ are a high-quality model and a poor model, respectively, $\boldsymbol{y}$ is the condition to perform the guidance on, and $w>1$ is the guidance scale. Typically the same network $D_{\theta}$ is used for $D_{1}$ and $D_{0}$ with additional degradations applied to $D_{0}$, such as corrupted input, e.g., CFG sets $D_{1}=D_{\theta}(\boldsymbol{z}_{s}|\boldsymbol{z}_{t},\boldsymbol{y})$ and $D_{0}=D_{\theta}(\boldsymbol{z}_{s}|\boldsymbol{z}_{t},\emptyset)$. The idea of autoguidance is that the weaker version of the same model amplifies the errors, and thus emphasizing the output of $D_{1}$ over $D_{0}$ by setting $w>1$ eliminates these errors.